OpenStudy (anonymous):

Solve the equation : v3x-5 = 4 v stands for sqrt A. 7 B. 3 C. 1/3 D. 49

OpenStudy (anonymous):

\[\sqrt{3x-5} = 4\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{3x-5} = 4\] \[\Large \left(\sqrt{3x-5}\right)^2 = 4^2\] \[\Large 3x-5 = 16\] \[\Large 3x-5+5 = 16+5\] \[\Large 3x = 21\] \[\Large \frac{3x}{3} = \frac{21}{3}\] \[\Large x = 7\]